OpenStudy (anonymous):

Find the function rule. x –2 –1 0 1 2 y 9 4 –1 –6 –11

myininaya (myininaya):

the x's are moving up by unit each time so look at the y's and see what the common difference is you have an arithmetic sequence

myininaya (myininaya):

but you don't know to know that term if you don't already

myininaya (myininaya):

you have a line find the slope and the y-intercept

OpenStudy (anonymous):

?

myininaya (myininaya):

what does ? mean

myininaya (myininaya):

like what do you have a question on?

myininaya (myininaya):

the slope or the y-intercept?

OpenStudy (anonymous):

how do i do it

myininaya (myininaya):

the slope formula is change of y/change of x the y-intercept is given (0,b) <---so what is b here?

myininaya (myininaya):

when x=0, y=? look at your table

OpenStudy (anonymous):

It might be helpful to think of it in terms of patterns. What's the pattern? What happens to y each time we go up by one?

OpenStudy (anonymous):

i still dont understand

myininaya (myininaya):

do you know how to read the table?

myininaya (myininaya):

it says the following points are on the line: (-2,9) (-1,4) (0,-1) (1,-6) (2,-11)

myininaya (myininaya):

I was asking you to tell me what y is when x=0

myininaya (myininaya):

the things I listed above are in the form (x,y)

myininaya (myininaya):

do you identify the ordered pair that has 0 as its first coordinate ?

myininaya (myininaya):

ok let's go back to the first thing I mentioned do you notice a common difference amongst the y values given like compare 9-4=? 4-(-1)=? -1-(-6)=? (-6)-(-11)=?

OpenStudy (anonymous):

5

myininaya (myininaya):

right that is the slope number y=5x+b now you just need to say what b is which is the y-intercept are you still not able to determine which pair (x,y) in the list below (-2,9) (-1,4) (0,-1) (1,-6) (2,-11) has 0 as the first coordinate?

myininaya (myininaya):

and should be y=-5x+b since the y values are decreasing

myininaya (myininaya):

but anyways look at the list of coordinates I just gave you it should be easy to tell which is of the form (0,b)

OpenStudy (anonymous):

is it 0 -1

myininaya (myininaya):

yes when x=0 y=-1

myininaya (myininaya):

So b=-1

myininaya (myininaya):

you have your function rule now

OpenStudy (anonymous):

Thank you so much!!!!!

myininaya (myininaya):

by the way can you tell me what it is just to be sure you know what it is

OpenStudy (anonymous):

is this what it should it look like this y=-5x+ -1

myininaya (myininaya):

that is right or you can write it as y=-5x-1

myininaya (myininaya):

either way

OpenStudy (anonymous):

thx